Combustion of an 8.23 mg sample of a compound gave 9.63 mg of carbon dioxide and 3.94 mg water. Analysis of another 5.32 mg sample of the same compound by the Carius method gave 13.49 mg of AgCl. 


1
Expert's answer
2022-02-23T01:04:05-0500

1) Find mass and moles of carbon

"n(C)= \frac{9.62\times 10^{-3}}{44.01} =0.000219 moles"

"m(C) =n\times M = 0.000219\times 12.01 = 0.00263 g =2.63 mg"


2) Find mass and moles of hydrogen

"n(H) = n(H_2O)\times 2 =\frac{3.94\times10^{-3}}{18.02} \times 2 = 0.000437 moles"

"m(H) = n\times M = 0.000437\times 1 = 0.437mg"


3) Find mass percent of chlorine in a sample

a) find mass of chlorine if silver nitrate:

"n(AgCl )= \frac{m}{M} = \frac{13.49\times 10^{-3}}{143.32} = 0.0000941 mol"

"n(Cl) = n(AgCl) = 0.0000941 mol"

"m(Cl) = 0.0000941\times35.45 = 0.00334 g = 3.34 mg"


b) find mass percent of chlorine in silver nitrate

"mass\% (Cl)= \frac{3.34}{5.32} =0.627"


4) Find mass and moles of chlorine in the first sample

"m(Cl) = m_{sample}\times mass\%(Cl) = 8.23\times 0.627 = 5.16 mg"

"n(Cl) = \frac{m}{M} = \frac{5.16\times10^{-3}}{35.45} = 0.000146 mol"


5) check up if there is an oxygen in a sample:

"8.23 = m(O) + m(C) + m(H) + m(Cl)"

"8.23 = m(O) +2.63 + 0.437 + 5.16"

"m(O) = 0.003 mg" (negligible)


6) find mole ratio

"n(C):n(H):n(Cl) = 0.000219:0.000438:0.000146 \n= 1.5:3:1 = 3:6:2"

The formula is "C_3H_6Cl_2"


7) Percent composition

mass percent "\% (c) = \frac{2.63}{8.23}\times 100\% = 32.0\%"

mass percent "\% (H)= \frac{0.437}{8.23} \times 100\% = 5.31 \%"

mass percent "\% =\frac{5.16}{8.23} \times 100\% = 62.7\%"